OpenStudy (jinx.exe):

Cassidy purchases a soccer ball that is regularly priced at $27.50. It is on sale for 30% off. What is the sale price of the soccer ball?

OpenStudy (ajanijones):

, Answer: Converting percentage to a decimal: 30% = 30 / 100 = 0.3 What is 30% of 27.5 = 30% * 27.5 = 0.3 * 27.5 = 8.25 Hope I helped :)